OpenStudy (anonymous):

Factor: 100x to the power of 4 - 81 A. x^2 (10x+9) (10x -9) B. (25x^2+9)(25x^2-9) C. (10x^2+9)(10x^2-9) D. (4x^2+9)(5x+3)(5x-3)

OpenStudy (anonymous):

\[100x^4-81\] start by factoring it as the difference of two squares \[a^2-b^2=(a+b)(a-b)\] with \[a=10x^2,b=9\]

OpenStudy (anonymous):

then?

OpenStudy (anonymous):

@AccessDenied

OpenStudy (accessdenied):

What do you have so far after doing that factor?

OpenStudy (anonymous):

i don't understand how to factor

OpenStudy (accessdenied):

This is a general rule of factoring, called "difference of squares" \( a^2 - b^2 = (a + b)(a - b) \) Have you seen this particular form used before?

OpenStudy (anonymous):

nope

OpenStudy (accessdenied):

To show that it is true, you could work backwards: \( \begin{align} ( a + b ) ( a - b) &= a (a - b) + b (a - b) \quad & \text{distribute} \\ & =a^2 - ab + ab - b^2 & \text{distribute again}\\ & = a^2 - b^2 & \text{cancel like terms}\end{align} \) I am not certain how to proceed if we don't have this form clear. \( a^2 - b^2 = (a + b)(a - b) \)

OpenStudy (accessdenied):

Some common examples: \( \color{red}{x^2 - 1} = \color{gray}{(x)^2 - (1)^2} =\color{maroon}{ (x + 1)(x - 1)} \) \( x^2 - 100 = \color{gray}{(x)^2 - (10)^2} = (x + 10)(x - 10) \) \( 4x^2 - 9 = \color{gray}{(2x)^2 - (3)^2} = (2x + 3)(2x - 3) \) \( x^4 - 1 = \color{gray}{(x^2)^2 - 1^2} = (x^2 + 1)\color{red}{(x^2 - 1)} = (x^2 + 1)\color{maroon}{(x + 1)(x - 1)} \) Back to the original question: Because with your original expression, you can rewrite as a difference of squares: \( 100x^4 - 81 \) = \( (\underbrace{10 x^2}_{a} )^2 - (\underbrace{9}_{b})^2 \) This form is met: \(a^2 - b^2 = (a + b)(a - b) \), where \(a = 10x^2\) and \(b = 9\): = \( (\underbrace{10x^2}_{a} + \underbrace{9}_{b})(\underbrace{10x^2}_{a} - \underbrace{9}_b) \)

OpenStudy (anonymous):

mhm

OpenStudy (accessdenied):

It really just comes down to memorizing that \(a^2 - b^2 = (a + b)(a - b) \). Most often, a binomial with perfect squares as terms and subtraction between the two numbers can be dealt with in this way... Sorry, I can't think of a better explanation at the moment if you've never used it. :(

OpenStudy (anonymous):

can you help me answer it

OpenStudy (anonymous):

@AccessDenied